OpenStudy (brucebaner):

What is the solution to the system of equations? x+y=5 x-2y=2 ([________]),([________[)

OpenStudy (brucebaner):

@KamiBug

OpenStudy (brucebaner):

can you help me

OpenStudy (brucebaner):

can you help me please :(

OpenStudy (kamibug):

You can solve the first equation for x and then plug it's value into the second equation to solve that one for y. Once you know what y is, plug it's value into the first equation to find x. So subtract y from both sides in x + y = 5 and you get x = 5 - y Now plug 5 - y for x into the second equation. (5 - y) - 2y = 2 Can you solve that for y for me? :)

OpenStudy (brucebaner):

@hlee13